PC Yaw Inc. approached Rudy and offered to lend him a capital for his 

computer business amounting to ₱300,000.00 to be paid in 3 years with 12% 

monthly interest. How much would be is his monthly payment?


1
Expert's answer
2022-01-09T17:01:22-0500

Use the formula:

"P = R\cdot\dfrac{1-(1+i/n)^{-nt}}{i/n}"

In finding "R," Regular (monthly) payment, we use: 


"R=P\cdot\dfrac{i/n}{1-(1+i/n)^{-nt}}"

Given "P=300000.00, i=0.12, n=12, t=3"


"R=300000\cdot\dfrac{0.12/12}{1-(1+0.12/12)^{-12(3)}}"

"R=9964.29"

His monthly payment would be "\u20b19964.29."